# WEATHER-5K: A Large-scale Global Station Weather Dataset Towards Comprehensive Time-series Forecasting Benchmark

### v1

@@ Appendix A Dataset Documentation
We organize the dataset documentation based on the template of datasheets for datasets [54].
A.1 Motivation
@@ A.2 Composition
Our dataset is collected and further processed using data sourced from the National Centers for Environmental Information (NCEI), specifically the Integrated Surface Database (ISD),  [^3] https://www.ncei.noaa.gov/products/land-based-station/integrated-surface-database. Although the ISD contains records from over $20,000$ stations spanning several decades, not all stations are suitable for machine learning applications. For instance, some stations are no longer operational, many do not report data on an hourly basis, and numerous stations have missing values for critical weather elements. To get a high-quality weather station dataset, a meticulous selection process was conducted to include only long-term, hourly reporting stations that are currently operational and provide essential observations such as temperature, dew point temperature, wind, and sea level pressure. After that, we use the process procedure detailed in Section 3 to make the final WEATHER-5K, which is in the principle of applicability for time-series forecasting research.
@@
Yes, the errors and noise in the dataset arise from two main sources. Firstly, the use of meteorological automatic stations introduces a certain degree of observational error, particularly in the measurement of wind speed and direction, which are relatively difficult to measure accurately. Secondly, in our data processing efforts to ensure the completeness of the dataset, we have employed interpolation operations, which can introduce some additional error. However, the proportion of error introduced by the interpolation is relatively small.
@@ A.3 Collection Process
A.4 Preprocessing/cleaning/labeling
@@
The raw data are available in the NCEI. The link is: https://www.ncei.noaa.gov/products/land-based-station/integrated-surface-database. To get the preprocessed data, you can run the ‘weather_station_api.py’ in our released repository: https://github.com/taohan10200/WEATHER-5K.
@@ A.5 Uses
We believe that our dataset will not encounter usage limit.
@@ A.6 Distribution
The instructions for building WEATHER-5K are available at: https://github.com/taohan10200/WEATHER-5K. The dataset does not have a digital object identifier currently.
@@
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No.
A.7 Maintenance
@@
Appendix B More Dataset Information
@@
Informer [34] Autoformer [13] Pyraformer [35] FEDformer [15] DLinear [40]
Train Time (Hours) 21$\sim$22 36$\sim$40 20$\sim$21 38$\sim$40 1.0$\sim$1.5
@@
PatchTST[37] Corrformer [12] Mamba [41] iTransformer [16]
Train Time (Hours) 7$\sim$8 144$\sim$168 3$\sim$4 3$\sim$4
GPU Memory 22,512 46,486 11,406 45,672
Appendix C More Experimental Results
@@
Visualization results. In Figures 7  8  9  10  11  12  13, we plot visualization results to showcase the performance of various time-series forecasting methods, including Pyraformer, FEDformer, DLinear, PatchTST, Mamba, iTransformer, and Corrformer. These visualizations provide a comparative analysis of how each of these different forecasting approaches performs on the time-series data.
@@
Appendix D Live Weather Demon

### v2

@@ Appendix B Dataset Documentation
We organize the dataset documentation based on the template of datasheets for datasets Gebru et al. (2021).
B.1 Motivation
@@ B.2 Composition
Our dataset is collected and further processed using data sourced from the National Centers for Environmental Information (NCEI), specifically the Integrated Surface Database (ISD), 111www.ncei.noaa.gov/products/land-based-station/integrated-surface-database. Although the ISD contains records from over $20,000$ stations spanning several decades, not all stations are suitable for machine learning applications. For instance, some stations are no longer operational, many do not report data on an hourly basis, and numerous stations have missing values for critical weather elements. To get a high-quality weather station dataset, a meticulous selection process was conducted to include only long-term, hourly reporting stations that are currently operational and provide essential observations such as temperature, dew point temperature, wind, and sea level pressure. After that, we use the process procedure detailed in Section 3 to make the final WEATHER-5K, which is in the principle of applicability for time-series forecasting research.
@@
Yes, the errors and noise in the dataset arise from two main sources. Firstly, the use of meteorological automatic stations introduces a certain degree of observational error, particularly in the measurement of wind speed and direction, which are relatively difficult to measure accurately. Secondly, in our data processing efforts to ensure the completeness of the dataset, we have employed interpolation operations, which can introduce some additional errors. However, the proportion of error introduced by the interpolation is relatively small.
@@ B.3 Collection Process
B.4 Preprocessing/cleaning/labeling
@@
The raw data are available in the NCEI. The link is: www.ncei.noaa.gov/products/land-based-station/integrated-surface-database. To get the preprocessed data, you can run the ‘weather_station_api.py’ in our released repository github.com/taohan10200/WEATHER-5K.
@@ B.5 Uses
We believe that our dataset will not encounter a usage limit.
@@ B.6 Distribution
The instructions for building WEATHER-5K will be available in the released code. The dataset does not have a digital object identifier currently.
@@
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation. No.
B.7 Maintenance
@@
Appendix C More Dataset Analysis
@@
Compare ISD with MADIS Data Source. Overall, as shown in Table 4, ISD is more diverse and offers broader coverage. Specifically, the surface data in MADIS mainly includes METER  [^2] https://madis.ncep.noaa.gov/madis_metar.shtml and Mesonet  [^3] https://madis.ncep.noaa.gov/madis_mesonet.shtml. Its reports primarily come from the U.S. while ISD collects surface weather data from more than $35,000$ stations worldwide. Additionally, ISD spans a longer period from ‘1901 to the present‘ and is fully public for users. In future research, we believe including MADIS for station-based weather forecasting will further enhance this field.
Table 4: Differences between ISD (database of WEATHER-5K) and MADIS
Dataset Availability Data Source Time Coverage
MADIS (METER and Mesonet) Restricted ASOS, AWOS, Airport Reports, CWOP, FAWH, GPSMet, KSDOT, RAWS, UDFCD, GLDNWS, IADOT, INTERNET 2001-present Primarily in U.S.
ISD Fully public More than 100 original data sources 1901-present 35,000 global stations
@@
Informer Autoformer Pyraformer FEDformer DLinear
Training Time (Hours) 21$\sim$22 36$\sim$40 20 $\sim$ 21 38 $\sim$ 40 1.0 $\sim$ 1.5
@@
Parameters (M) 11.32 10.53 7.54 16.29 0.01
PatchTST Corrformer Mamba iTransformer
Training Time (Hours) 7$\sim$8 144$\sim$168 3$\sim$4 3$\sim$4
GPU Memory (MiB) 22,512 46,486 11,406 45,672
Parameters (M) 6.65 666.12 0.12M 4.8
Appendix D More Experimental Results
@@
Visualization results. In Figures 8  9  10  11  12  13  14, we have plot visualization results to showcase the performance of various time-series forecasting methods, including Pyraformer, FEDformer, DLinear, PatchTST, Mamba, iTransformer, and Corrformer. These visualizations provide a comparative analysis of how each of these different forecasting approaches performs on the time-series data.
@@
Appendix E Live Weather Demon